OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10?

OpenStudy (hba):

m=-b/a

OpenStudy (hba):

x/2+y/5=1

OpenStudy (hba):

a=2,b=5

OpenStudy (hba):

m=-5/2

OpenStudy (anonymous):

ty very much

OpenStudy (hba):

welcome